OpenStudy (anonymous):

Selena orders CDs for a music shop. She just ordered 8 CDs from each of 11 pop artists and 5 CDs from each of 8 jazz artists. How many CDs did she order altogether?

OpenStudy (anonymous):

\[(8\times11) + (5\times8)=88+40=128\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problem!